# Anisotropy can make a moving active fluid membrane rough or crumpled

### v1

@@ C.4 One-loop corrections to $\nu_{ij}$
$-\lambda_{ij}\lambda_{mn}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}q_{m}k_{n}\langle|h_{{\bf q},\Omega}|^{2}\rangle\langle h_{{\bf k-q},-\Omega}\hat{h}_{{\bf-k+q},\Omega}\rangle.$
After symmetrizing, i.e., ${\bf q}\rightarrow{{\bf q}+{{\bf k}}/{2}}$, and performing the $\Omega$-integral, we are left with two integrals that include corrections for $\nu$ and $\overline{\nu}$. These integrals are
$\mathcal{I}_{1}=\frac{\lambda_{ij}\lambda_{mn}k_{n}k_{m}D}{4}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{q_{i}q_{j}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}},$ (32)
@@
$\mathcal{I}_{2}=-\frac{\lambda_{ij}\lambda_{mn}k_{n}D}{2}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{q_{i}q_{j}q_{m}\bigl{[}\nu(k_{x}q_{x}+k_{y}q_{y})+\overline{\nu}(k_{x}q_{y}+k_{y}q_{x})\bigl{]}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{3}}.$ (33)
@@
Corrections to $\nu$ are given by the coefficient of $k_{x}^{2}$ (or $k_{y}^{2}$). Considering the coefficient of $k_{x}^{2}$, $\lambda_{ij}\lambda_{mn}k_{n}k_{m}q_{i}q_{j}=k_{x}^{2}\lambda(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})$. Substituting this in Eq. (32), the correction to $\nu$ coming from $\mathcal{I}_{1}$ is
$\frac{D\lambda k_{x}^{2}}{4}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}}=\frac{D\lambda k_{x}^{2}}{4\nu^{2}}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{(q_{x}-a_{+})^{2}(q_{x}-a_{-})^{2}}.$ (34)
To evaluate the $q_{x}$-integral, considering a closed contour in the upper half plane the relevant pole $q_{x}=a_{+}=q_{y}(-\alpha+ip)$, which is a second order pole. The residue at $q_{x}=a_{+}$ is
$\frac{d}{dq_{x}}\Biggl{[}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{(q_{x}-a_{-})^{2}}\Biggl{]}_{q_{x}=a_{+}}=-i\frac{\lambda-\overline{\lambda}\alpha}{2p^{3}q_{y}}.$
Multiplying by $2\pi i$ and writing $p$ in terms of $\alpha$, Eq. (34) becomes
$k_{x}^{2}\frac{\lambda D(\lambda-\overline{\lambda}\alpha)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{(2\pi)q_{y}}.$ (35)
Equation (35) gives the correction to $\nu$ coming from $\mathcal{I}_{1}$. Similarly keeping only order of $k_{x}^{2}$, Eq. (33) becomes
$-k_{x}^{2}\frac{D}{2\nu^{3}}\int_{q_{x},q_{y}}\biggr{[}\frac{\lambda^{2}\nu q_{x}^{4}+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})q_{x}^{2}q_{y}^{2}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}+\frac{(3\lambda\overline{\lambda}\nu+\lambda^{2}\overline{\nu})q_{x}^{3}q_{y}+(\lambda^{2}\overline{\nu}+2\overline{\lambda}^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)q_{x}q_{y}^{3}+\lambda\overline{\lambda}\overline{\nu}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}\biggr{]}.$ (36)
This integral is similar to the integral in Eq. (27). However, $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ are now different with values:
$\mathcal{A}=\lambda^{2}\nu,\mathcal{B}=(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu}),\mathcal{C}=3\lambda\overline{\lambda}\nu+\lambda^{2}\overline{\nu},\mathcal{D}=\lambda^{2}\overline{\nu}+2\overline{\lambda}^{2}\overline{\nu}+\lambda\overline{\lambda}\nu,\mathcal{E}=\lambda\overline{\lambda}\overline{\nu},$
Comparing with Eq. (29) and substituting these values of $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ equation (36) can be written as
$-k_{x}^{2}\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\Biggl{[}(\lambda^{2}\nu+\lambda\overline{\lambda}\overline{\nu})-(2\lambda^{2}\overline{\nu}+4\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu})\alpha+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})\Bigr{(}\frac{1+2\alpha^{2}}{3}\Bigr{)}\Biggl{]}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (37)
Equation (37) gives the correction to $\nu$ coming from $\mathcal{I}_{2}$. Adding Eqs. (35) and (37) and extracting the coefficient of $k_{x}^{2}$, we get the total correction to $\nu$ at the one-loop order as
$-\frac{3D}{32\nu^{3}{(1-\alpha^{2})^{\frac{5}{2}}}}\frac{\ln b}{2\pi}\Biggl{[}(\lambda^{2}\nu+\lambda\overline{\lambda}\overline{\nu})-(2\lambda^{2}\overline{\nu}+4\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu})\alpha+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})\Bigr{(}\frac{1+2\alpha^{2}}{3}\Bigr{)}\Biggl{]}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{D}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\Biggl{[}\lambda^{2}-\lambda\overline{\lambda}\alpha\Biggl{]}.$ (38)
Substituting $\alpha={\overline{\nu}}/{\nu}$ in above equation and after some algebra the above expression reduces to
$\frac{\nu D}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{5}{2}}}\frac{\ln b}{2\pi}\Biggl{[}(\nu^{2}-\overline{\nu}^{2})\Big{(}\frac{\lambda\overline{\lambda}\overline{\nu}}{\nu}-\overline{\lambda}^{2}\Big{)}\Biggl{]}.$ (39)
To get the corrections to $\overline{\nu}$, we need to find coefficient of $2k_{x}k_{y}$ from Eqs. (32) and (33). Therefore coonsidering only the $k_{x}k_{y}$-term, $\lambda_{ij}\lambda_{mn}k_{n}k_{m}q_{i}q_{j}\rightarrow 2k_{x}k_{y}\overline{\lambda}\bigl{[}\lambda(q_{x}^{2}+q_{y}^{2})+2\overline{\lambda}q_{x}q_{y}\bigl{]}$. Substituting this in Eq. (32) we get
$\frac{D\overline{\lambda}(2k_{x}k_{y})}{4}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}}.$ (40)
The integral in Eq. (40) is same as the integral in Eq. (34). Comparing with Eq. (35) we can write Eq. (40) as
$(2k_{x}k_{y})\frac{D\overline{\lambda}(\lambda-\overline{\lambda}a)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (41)
Equation (41) gives the correction to $\overline{\nu}$ that comes from $\mathcal{I}_{1}$. Now,
$\lambda_{ij}\lambda_{mn}k_{n}q_{i}q_{j}q_{m}=k_{x}(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})(\lambda q_{x}+\overline{\lambda}q_{y})+k_{y}(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})(\lambda q_{y}+\overline{\lambda}q_{x}).$
Substituting this in Eq. (33) and extracting the coefficient of $2k_{x}k_{y}$ we get
$-(2k_{x}k_{y})\frac{D}{2\nu^{3}}\int_{q_{x},q_{y}}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}.$ (42)
Notice that the integral in Eq. (42) has the same form as the integral in Eq. (28), but now $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ are given by $\mathcal{A}=\mathcal{E}=\bigr{(}\frac{\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu}{2}\bigr{)},\mathcal{B}=(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu}),\mathcal{C}=\mathcal{D}=(\lambda^{2}\nu+\overline{\lambda}^{2}\nu+2\lambda\overline{\lambda}\overline{\nu})$. Thus comparing with Eq. (29) we write Eq. (42) as
$-(2k_{x}k_{y})\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)+(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu})\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-2(\lambda^{2}\nu+\overline{\lambda}^{2}\nu+2\lambda\overline{\lambda}\overline{\nu})\alpha\biggr{]}.$ (43)
Equation (43) gives the correction to $\overline{\nu}$ that comes from $\mathcal{I}_{2}$. Thus adding Eq. (41) and (43) and extracting the coefficient of $2k_{x}k_{y}$, we get total correction to $\overline{\nu}$ given by
$-\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)+(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu})\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-(2\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+4\lambda\overline{\lambda}\overline{\nu})\alpha\biggr{]}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{D(\overline{\lambda}\lambda-\overline{\lambda}^{2}\alpha)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (44)
Substituting $\alpha={\overline{\nu}}/{\nu}$ in the above expression and after some algebraic simplification the above contribution becomes
$\frac{\lambda^{2}\overline{\nu}D}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\Biggl{[}1-\frac{\overline{\lambda}\nu}{\lambda\overline{\nu}}\Biggl{]}.$ (45)
C.5 Nonrenormalization of $\lambda$ and $\overline{\lambda}$
Figure 9: One-loop Feynman diagrams giving the corrections to $\lambda$ and $\overline{\lambda}$.
@@ C.5 Nonrenormalization of $\lambda$ and $\overline{\lambda}$
The left diagram in Fig. 9 has a symmetry factor 24 and includes corrections to both $\lambda$ and $\overline{\lambda}$. We find
$-\lambda_{ij}\lambda_{mn}\lambda_{rs}D\int_{q,\Omega}\frac{(q_{i}q_{j})(-k_{1m}q_{n})(q_{r}k_{2s})}{(-i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})(i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})\Bigl{(}\Omega^{2}+(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{)}}$
$=\frac{\lambda_{ij}\lambda_{mn}\lambda_{rs}D}{4}\int_{q}\frac{(q_{i}q_{j})(k_{1m}q_{n})(q_{r}k_{2s})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{3}}.$ (46)
@@
$-\lambda_{ij}\lambda_{mn}\lambda_{rs}(2D)\int_{q,\Omega}\frac{(q_{i}q_{j})(-k_{1m}q_{n})(-q_{r}k_{2s})}{(-i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{(}\Omega^{2}+(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{)}}$
$=-\frac{\lambda_{ij}\lambda_{mn}\lambda_{rs}D}{4}\int_{q}\frac{(q_{i}q_{j})(k_{1m}q_{n})(q_{r}k_{2s})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{3}}.$ (47)
Equations (46) and (47), when added, give zero contribution to the correction of $\lambda$ and $\overline{\lambda}$. These are consistent with the Galilean invariance of Eq. (1) in the main text.
C.6 RG recursion relations
We integrate out the dynamical fields with wavevectors $k_{y}$ in the range ($\Lambda/b\leq k_{y}\leq\Lambda$ and $-\Lambda\leq k_{y}\leq-\Lambda/b$) in the action functional (20). Now we rescale space, time and fields to raise the wavevector cutoff of $k_{y}$ to ±$\pm\Lambda$. We rescale ${\bf k}\to{\bf k}^{{}^{\prime}}=b{\bf k}\implies{\bf x}\to{\bf x}^{{}^{\prime}}=\frac{{\bf x}}{b}$ similarly, $\omega\to\omega^{{}^{\prime}}=b^{z}\omega\implies t{{}^{\prime}}=\frac{t}{b^{z}}$. Furthermore, the fields are rescaled as

### v2

@@ III.4 One-loop corrections to $\nu_{ij}$
$-2D\lambda^{2}\int_{{\bf q},\Omega}\frac{q_{i}(k-q)_{i}q_{j}k_{j}}{\big[i\Omega+\nu_{x}(k-q)_{x}^{2}+\nu_{y}(k-q)_{y}^{2}+2\overline{\nu}(k-q)_{x}(k-q)_{y}\big]\big[\Omega^{2}+(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{2}\big]}.$
After symmetrizing, i.e., ${\bf q}\rightarrow{{\bf q}+{{\bf k}}/{2}}$, and performing the $\Omega$-integral, we are left with two integrals that include corrections for $\nu_{x}$, $\nu_{y}$ and $\overline{\nu}$. These integrals are
$\mathcal{I}_{1}=(k_{x}^{2}+k_{y}^{2})\frac{\lambda^{2}D}{16\nu_{x}^{2}p^{3}}(1+\beta)\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$ (31)
@@
$\mathcal{I}_{2}=-2D\lambda^{2}\int_{q_{x},q_{y}}\frac{(q_{x}k_{x}+q_{y}k_{y})q^{2}}{4(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{3}}\big[k_{x}(\nu_{x}q_{x}+\overline{\nu}q_{y})+k_{y}(\nu_{y}q_{y}+\overline{\nu}q_{x})\big].$ (32)
@@
Corrections to $\nu_{x}$ ($\nu_{y}$) are given by the coefficient of $k_{x}^{2}$ ($k_{y}^{2}$). Considering the coefficient of $k_{x}^{2}$, the correction to $\nu_{x}$ coming from $\mathcal{I}_{2}$ is
$-k_{x}^{2}\frac{\lambda^{2}D}{2}\int_{q_{x},q_{y}}\frac{\nu_{x}q_{x}^{4}+\nu_{x}q_{x}^{2}q_{y}^{2}+\overline{\nu}q_{x}^{3}q_{y}+\overline{\nu}q_{y}^{3}q_{x}}{(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{3}},$ (33)
$=-k_{x}^{2}\frac{\lambda^{2}D}{2}\big[\nu_{x}I_{4}+\nu_{x}I_{5}+\overline{\nu}I_{6}+\overline{\nu}I_{7}\big],$ (34)
$=-k_{x}^{2}\frac{\lambda^{2}D}{2}\frac{3}{16\nu_{x}^{2}p^{3}}\bigl[\beta+1/3\bigl]\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (35)
Thus collecting coefficient of $k_{x}^{2}$ from $\mathcal{I}_{1}$ and adding it with above equation we obtain correction to $\nu_{x}$,
$\frac{\lambda^{2}D}{16\nu_{x}^{2}p^{3}}\frac{\ln b}{2\pi}\bigg[\frac{1}{2}-\frac{\beta}{2}\bigg].$ (36)
To compute the correction to $\nu_{y}$, we extract the coefficient of $k_{y}^{2}$ from integral $\mathcal{I}_{2}$,
$-k_{y}^{2}(2D\lambda^{2})\int_{q_{x},q_{y}}\frac{(q_{x}^{2}q_{y}+q_{y}^{3})(\nu_{y}q_{y}+\overline{\nu}q_{x})}{4(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{3}},$ (37)
$=-k_{y}^{2}\frac{\lambda^{2}D}{2}\big[\nu_{y}I_{5}+\nu_{y}I_{8}+\overline{\nu}I_{6}+\overline{\nu}I_{7}\big],$ (38)
$=-k_{y}^{2}\frac{\lambda^{2}D}{16\nu_{x}^{2}p^{3}}\bigl[3/2+\beta/2\bigl]\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (39)
Again collecting coefficient of $k_{y}^{2}$ from $\mathcal{I}_{1}$ and adding it with above equation we obtain correction to $\nu_{y}$,
$\frac{\lambda^{2}D}{16\nu_{x}^{2}p^{3}}\frac{\ln b}{2\pi}\bigg[\frac{\beta}{2}-\frac{1}{2}\bigg].$ (40)
Next we collect coefficient of $2k_{x}k_{y}$ from $\mathcal{I}_{2}$ to obtain correction to $\overline{\nu}$ which is,
$-2k_{x}k_{y}\lambda^{2}D\int_{q_{x},q_{y}}\frac{q^{2}q_{x}(\nu_{y}q_{y}+\overline{\nu}q_{x})+q^{2}q_{y}(\nu_{x}q_{x}+\overline{\nu}q_{y})}{4(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{3}},$ (41)
$=-(2k_{x}k_{y})\frac{\lambda^{2}D}{4}\big[(\nu_{x}+\nu_{y})(I_{6}+I_{7})+\overline{\nu}(I_{4}+I_{8}+2I_{5})\big].$ (42)
After some algebra, and collecting coefficient of $2k_{x}k_{y}$ from above equation we obtain the correction to $\overline{\nu}$ which is $\frac{\lambda^{2}D\alpha}{16\nu_{x}^{2}p^{3}}\frac{\ln b}{2\pi}.$
III.5 Nonrenormalization of $\lambda$
Figure 8: One-loop Feynman diagrams giving the corrections to $\lambda$.
@@ III.5 Nonrenormalization of $\lambda$
The left diagram in Fig. 8 has a symmetry factor 24 and includes corrections to $\lambda$. We find
$-\lambda^{3}D\int_{{\bf q},\Omega}\frac{q^{2}(-k_{1m}q_{m})(q_{n}k_{2n})}{(-i\Omega+\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})(i\Omega+\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})\Bigl(\Omega^{2}+(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}\Bigl)},$
$=\frac{\lambda^{3}D}{4}\int_{\bf q}\frac{q^{2}(k_{1m}q_{m})(q_{n}k_{2n})}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}.$ (43)
@@
$-\lambda^{3}(2D)\int_{{\bf q},\Omega}\frac{q^{2}(-k_{1m}q_{m})(-q_{n}k_{2n})}{(-i\Omega+\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}\Bigl[\Omega^{2}+(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}\Bigl]},$
$=-\frac{\lambda^{3}D}{4}\int_{\bf q}\frac{q^{2}(k_{1m}q_{m})(q_{n}k_{2n})}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}.$ (44)
Equations (43) and (44), when added, give zero contribution to the correction of $\lambda$. These are consistent with the Galilean invariance of Eq. (1).
III.6 RG recursion relations
We integrated out the dynamical fields with wavevectors $k_{y}$ in the range ($\Lambda/b\leq k_{y}\leq\Lambda$ and $\Lambda\leq k_{y}\leq-\Lambda/b$) in the action functional (23). Now we rescale space, time and fields to raise the wavevector cutoff of $k_{y}$ to $\pm\Lambda$. We rescale ${\bf k}\to{\bf k}^{{}^{\prime}}=b{\bf k}\implies{\bf x}\to{\bf x}^{{}^{\prime}}=\frac{{\bf x}}{b}$ similarly, $\omega\to\omega^{{}^{\prime}}=b^{z}\omega\implies t{{}^{\prime}}=\frac{t}{b^{z}}$. Furthermore, the fields are rescaled as